# Collective emission and selective-radiance in atomic clouds and arrays coupled to a microring resonator

### v1

@@ IV Conclusion
We studied the spectrum and decay rates of a cloud of atoms trapped near the resonator. We explored how the two different types of excitation, using a short pulse (for the timed-Dicke state) or a long pulse (for the steady state), affect the decay dynamics of the system. We also studied how the ratio of the decay rate into the desired and undesired modes scales with the number of atoms.
We explored the implications of trapping atoms in sub-wavelength arrays along the resonator. We studied the effects of the high wavenumber excitation through the microring and explored potential impacts of disorder on subradiance in experimental implementations. This platform also holds the potential to form a ring of trapped atoms along the microring resonator to achieve a much better scaling on subradiance when compared to linear arrays of atoms.
@@ Appendix A Adiabatic Elimination of the cavity
The cavity has two sources of loss. It can intrinsically lose a photon with the rate $\kappa_{i}$. Cavity photons can also hop to the bus waveguide with coupling rate $\kappa_{e}$. This cavity decay is defined by the Lindblad operator,
@@
By enforcing that the cavity reaches a steady state at every time-step of the internal dynamic evolution ($\langle\dot{\hat{a}}\rangle=0$), we arrive at an expression for $\langle\hat{a}\rangle$ depending on the atomic excitations.
@@
Appendix B Orthogonality of the eigenmodes
The atom-cavity interaction matrix $\mathbf{G}_{c}$ is skew-symmetric because of the running-wave nature of the resonator. This results in purely imaginary eigenvalues, meaning none of the eigenmodes of the atom-cavity interaction has any energy shift. These eigenmodes also follow the familiar orthogonality relation $\vec{U}\cdot\vec{V}=\vec{U}^{\dagger}\vec{V}$.
On the other hand, the interaction matrix of the free space modes $\mathbf{G}_{f}$ is complex-symmetric. This means that $G^{f}_{ij}=G^{f}_{ji}$. This results in the eigenmodes having a different orthogonality relation $\langle\vec{U},\vec{V}\rangle=\vec{U}\vec{V}$.
@@ Appendix B Orthogonality of the eigenmodes
$\begin{split}\mathbf{M}\vec{R}_{\alpha}=\lambda_{\alpha}\vec{R}_{\alpha}\\ \vec{L}^{\dagger}_{\alpha}\mathbf{M}=\lambda_{\alpha}\vec{L}^{\dagger}_{\alpha},\end{split}$ (29)
@@
$\vec{L}^{\dagger}_{\alpha}\vec{R}_{\alpha^{\prime}}=\delta_{\alpha,\alpha^{\prime}}.$ (30)
@@
$\tilde{\Omega}_{\alpha}=\vec{L}^{\dagger}_{\alpha}\vec{\Omega}.$ (31)
@@
with the proportionality depending on the pulse area. These can be used with Eq. (20) to calculate the final state and the composition with respect to eigenmodes.

### v2

@@ V Conclusion
We studied the spectrum and decay rates of a cloud of atoms trapped near the resonator. We explored how the two different types of excitation, using a short pulse (for the timed-Dicke state) or a long pulse (for the steady state), affect the decay dynamics of the system. Although a suppression of the decay rate can be seen with the steady state, achieving selective radiance would be difficult in randomized atomic clouds.
Hence, we explored the implications of the proposed plan of trapping atoms in sub-wavelength arrays along the resonator. We show that this could afford the coherence necessary to achieve selective radiance. We studied the effects of the high-wavenumber excitation through the microring and explored potential impacts of disorder on subradiance in experimental implementations.
@@
The data plotted in the figures are openly available in Ref. Suresh et al. (2025).
@@ Appendix A Adiabatic Elimination of the cavity
This appendix briefly describes the process of adiabatically eliminating the cavity photons from the Hilbert space due to the separation of timescales.
Beyond the cavity interaction Hamiltonian given in Eq. (10), the cavity has two sources of loss. It can intrinsically lose a photon with the rate $\kappa_{i}$. Cavity photons can also hop to the bus waveguide with coupling rate $\kappa_{e}$. This cavity decay is defined by the Lindblad operator,
@@
Since the cavity decay timescales are two orders of magnitude faster than the relevant atomic excitation timescales, they can be adiabatically eliminated. By enforcing that the cavity reaches a steady state at every time-step of the internal dynamic evolution ($\langle\dot{\hat{a}}\rangle=0$), we arrive at an expression for $\langle\hat{a}\rangle$ depending on the atomic excitations.
@@
Appendix B Orthogonality of the eigenmodes and final states
This appendix discusses the orthogonality relations in the combined Green’s function matrix and demonstrates how they can be used to directly compute the final states. These unfamiliar orthogonality relations arise due to the different non-Hermitian properties of the relevant matrices.
The atom-cavity interaction matrix $\mathbf{G}_{c}$ is skew-symmetric because of the running-wave nature of the resonator. This results in purely imaginary eigenvalues, meaning none of the eigenmodes of the atom-cavity interaction has any energy shift. Any two eigenmodes $\vec{V}_{\alpha},\vec{V}_{\beta}$ follows the familiar orthogonality relation: $\vec{V}_{\alpha}^{\dagger}\vec{V}_{\beta}=\delta_{\alpha\beta}$, where $\delta_{\alpha\beta}$ is the Kronecker delta.
On the other hand, the interaction matrix of the free space modes $\mathbf{G}_{f}$ is complex-symmetric. This means that $G^{f}_{ij}=G^{f}_{ji}$. This results in the eigenmodes having a different orthogonality relation: $\vec{V}_{\alpha}^{T}\vec{V}_{\beta}=\delta_{\alpha\beta}$, where $T$ denotes a transpose without conjugation.
@@ Appendix B Orthogonality of the eigenmodes and final states
$\begin{split}\mathbf{M}\vec{R}_{\alpha}=\lambda_{\alpha}\vec{R}_{\alpha}\\ \vec{L}_{\alpha}^{T}\mathbf{M}=\lambda_{\alpha}\vec{L}_{\alpha}^{T},\end{split}$ (31)
@@
$\vec{L}_{\alpha}^{T}\vec{R}_{\beta}=\delta_{\alpha\beta}.$ (32)
@@
$\tilde{\Omega}_{\alpha}=\vec{L}_{\alpha}^{T}\vec{\Omega}.$ (33)
@@
with the proportionality depending on the pulse area. These can be used with Eq. (21) to calculate the final state and the composition with respect to the eigenmodes.